how many quarts of pure antifreeze must be added to 9 quarts of a 20% antifreeze solution to obtain a 30% antifreeze solution?
Answer by modi.kk91(20)

let pure antifreeze=100%
quarts=x

x*100%+9*20%=30%(X+9)
100x+180=30x+270

100x-30x=270-180
70x= 90

x=90/70